# Multi-View Projection for Unsupervised Domain Adaptation in 3D Semantic Segmentation

### v2

@@ IV-A1 Datasets
SemanticKITTI [3] (SK) and nuScenes [2] (NS) are two well-known large-scale datasets for AD, both providing and annotating Lidar 3D pointclouds. The procedure generally adopted in the domain adaptation literature [24, 25, 18, 19]
@@
is to train on the official training splits for both datasets, e.g. 700 scenes for nuScenes and sequences 00-07 and 09-10 for SemanticKITTI, and evaluate on the official validation splits, e.g. 150 scenes for nuScenes and sequences 08 for SemanticKITTI. We follow this procedure for comparing with the state of the art. Of the 700 nuScenes scenes, we use the 70 last scenes for validation. Similarly, we use sequence 10 of SemanticKITTI for validation. For the rare classes experiment IV-E2 and the ablation study IV-F, we use a subset of the first 1800 scans from sequence 08 of the SK dataset for evaluation, to reduce the computational cost.
Simulation-to-Real
@@
SynLidar [4] (SL) is a synthetic AD dataset captured with Unreal Engine 4. The sub-sampled version that is used in most of the literature [4, 20, 21, 22] is composed of 13 sequences, containing 19,865 scans with annotations for 32 classes. No official validation split is provided, therefore we follow [24] and use sequences 05 and 10 for validation. We do not evaluate on SynLidar, as it is only used as a source dataset.
IV-A2 Metrics
@@
We measure the 3D semantic segmentation performance using the Intersection over Union (IoU) metric for all the experiments. We report both the per-class IoU and the mean IoU over all classes (mIoU). We compute all the metrics using code provided by [26], to ensure a fair comparison. For all UDA experiments, the evaluation is done on the 3D model after training on the pseudo-labels provided by Seg_3D_by_PC2D . For the rare classes and ablation experiments, the evaluation is directly done on the labels generated by the pipeline.
IV-B Implementation details
There are many hyper-parameters in our method that allow to try many different approaches. We present the most important ones here, and how we set them for best performance. Detailed parameters used will be available on github111https://github.com/andrewcaunes/ia4markings.
To begin with, we prepare the aligned scenes from SemanticKITTI and nuScenes. We use the already existing scenes from NS and accumulate scans 60 by 60 to create SK scenes. The intensities are normalized by scan, following [12]. The RGB images are used to colorize the pointclouds in both datasets using 2D-to-3D projection and a depth buffer to apply occlusion. Note that SynLidar does not provide RGB images, we therefore do not use RGB colors for the Sim-to-Real setting, where we use only 3 modalities: intensities, depths and normals.
@@ IV-B Implementation details
NuScenes’ annotations are only available at 2Hz, while the scans are captured at 20Hz. To prepare for 2D segmentation label generation for PC2D datasets, we make the annotations denser by using a simple KNN method with to assign labels to each point in the dense scene.
@@
The PC2D datasets are generated with 400,000 samples for each modality and source dataset. The views are uniformly sampled over the training scenes. For inference, as mentioned in III-C, we use 250 images per camera pose type, thus 1,000 images per scene, per modality. This amounts to 4,000 images total per scene. The inference time for each scene is relatively long, ~ 2 minutes on a NVIDIA RTX 3070 GPU, thus why the pseudo-label generation and fine-tuning application would be favored for real-time applications.
All views are rendered at 1024512 resolution with a standard field of view..
@@
IV-B1 Class mappings
We evaluate two Real-to-Real settings, NSSK and SKNS, and one Sim-to-Real setting, SLSK. To compare with the state-of-the-art, we follow [19, 21] and use 10 classes common to NS and SK. For the SLSK scenario, we use the official SK19 class mapping as in [20, 21]. To denote the different versions of the SemanticKITTI dataset, we use the following notations: SK10, SK19. The exact class mappings used can be found on github††footnotemark: .
| SLSK19 (% IoU) | Car | Bicycle | Motorcycle | Truck | Other vehicle | Pedestrian | Bicyclist | Motorcyclist | Road | Parking | Sidewalk | Other ground | Building | Fence | Vegetation | Trunk | Terrain | Pole | Traffic sign | % mIoU | |
| Source-only | 50.8 | 7.6 | 11.8 | 1.4 | 3.2 | 15.8 | 34.8 | 3.9 | 30.0 | 4.5 | 34.5 | 0 | 27.0 | 15.8 | 62.9 | 24.4 | 43.0 | 18.9 | 6.4 | 20.9 | |
| CoSMix | [20] | 75.1 | 6.8 | 29.4 | 27.1 | 11.1 | 22.1 | 25.0 | 24.7 | 79.3 | 14.9 | 46.7 | 0.1 | 53.4 | 13 | 67.7 | 31.4 | 32.1 | 37.9 | 13.4 | 32.2 |
| SALUDA | [21] | 67.0 | 7.7 | 14.4 | 1.3 | 5.2 | 24.1 | 52.6 | 2.7 | 52.5 | 10.5 | 44.1 | 0.4 | 51.8 | 13.6 | 69.7 | 40.5 | 56.5 | 45.0 | 14.3 | 30.2 |
| Unimix | [22] | 80.3 | 6.1 | 32.5 | 29.2 | 10.6 | 23.7 | 30 | 24.5 | 62.2 | 15.9 | 46.5 | 0.1 | 60.9 | 15.8 | 70.7 | 34.9 | 41.2 | 38.6 | 17.1 | 33.7 |
| Seg_3D_by_PC2D | (ours) | 83.5 | 0.9 | 6 | 10.4 | 3.0 | 19.8 | 13.5 | 0.2 | 80.2 | 0.0 | 51.6 | 0 | 78.8 | 16.7 | 51.9 | 32.0 | 36.8 | 40.6 | 10.1 | 28.2 |
IV-C Baselines
@@
To demonstrate the capabilities of our method, we compare it to the state-of-the-art methods in UDA for 3D semantic segmentation. We use the results reported by the authors for each method, while ensuring a fair comparison by using the same dataset splits,class mappings, metrics and model architectures. For Real-to-Real, in the NSSK10 setting, we find the best performing methods to date are Lidar-UDA [19], T-UDA [18], CoSMix [20] and SALUDA [21]. In the SK10NS setting, we compare with Lidar-UDA [19] and T-UDA [18] as CoSMix [20] and SALUDA [21] do not provide results for this setting. Note that T-UDA [18] also reports values for the manmade class, not reported by the other methods. We therefore do not include it and re-compute the accordingly. For Sim-to-Real, in the SLSK19 setting, we compare with SALUDA [21], CoSMix [20] and UniMix [22].
IV-D Unsupervised Domain Adaptation Results
IV-D1 Real-to-Real setting
@@
Results for the NSSK10 setting can be found in Table I. This UDA setting is particularly challenging as the source (NS) uses a 32-beam Lidar, less dense than the 64-beam Lidar used in the target (SK). Seg_3D_by_PC2D is able to outperform the state-of-the-art in mIoU by 1.75 Interestingly, our method reaches state-of-the-art performance even for dynamic classes such as car and motorcycle. This demonstrates that the model learns to match ‘shadow effect’-like features with the corresponding classes. Also reported in Table I is the supervised performance. This corresponds to the evaluation of the pseudo-labels outputed by the Seg_3D_by_PC2D pipeline using 2D models trained on the target dataset directly. It is interesting to note that in this ‘supervised’ setting, our method reaches relatively low performance compared to typical supervised methods (e.g. 65.3 mIoU in [21]), while still prevailing in UDA. We believe this is because the fully 2D reliance of the approach allows to bridge the domain adaptation gap successfully, while also hindering the potential performance by introducing errors in the projection and voting phase of the pipeline. In addition, it seems that the fine tuning of the pretrained model on pseudo-labels improves adaptation significantly, and the score of the pseudo-labels themselves does not act as a lower bound, indicating that there may exist a synergy between the source labels and target pseudo-labels that allows to reduce the domain gap.
@@
Results for the SK10NS setting can be found in Table II. In this setting, Seg_3D_by_PC2D is even more dominant, with a 5.13 over state-of-the-art. In particular, the method is superior on all static classes, which can be explained by the higher difficulty to segment dynamic classes displaying some ‘shadow effect’. We notice relatively lower performance on small classes, such as bicycle and pedestrian, which intuitively is linked to the higher rate of precision errors in the projections.
IV-D2 Simulation-to-Real setting
@@
Table III reports the results for the SLSK19 setting. Overall, our method achieves lower mIoU compared to prior works such as CoSMix [20], SALUDA [21], and UniMix [22]. However, it consistently performs strongly on several static classes, notably, road (80.24), sidewalk (51.55), and building (78.8), and some dynamic classes such as car (83.51), where it reaches state-of-the-art performance. This suggests that the projection-based pseudo-labeling is particularly effective on large, well-structured classes but less robust on fine-grained or rare classes such as bicycle, motorcycle, and motorcyclist. While the global mIoU is lower (28.21), the results highlight that our approach can still surpass prior methods on key static categories, underlining its strength for Sim-to-Real adaptation in structured environments
IV-E Application to rare classes
@@
A very common real world scenario in 3D semantic segmentation is to have access to 3D annotations for some classes in a source domain, but not the desired target classes. However, 2D annotations are more generally available as they are less costly. We show that our method can be used in such cases: by leveraging the 3D annotations and our PC2D dataset generation pipeline, a 2D model is pretrained in the source domain. Then, using the 2D annotations, the model is fine-tuned on the target classes. This is an improvement over [12], which only uses 2D annotations. The pretraining on 3D annotations allows to significantly bridge the domain gap that exists between camera images and rendered images.
IV-E1 Experimental setup
For this experiment, we need 3 datasets:
@@ IV-E1 Experimental setup
A pretraining 3D source dataset, with 3D annotations for any classes. We use nuScenes.
@@
A target 3D dataset. We use SemanticKITTI, as it conveniently provides 3D annotations for our target class, traffic sign, which will allow to perform a quantitative evaluation.
We compare Seg_3D_by_PC2D with seg_3D_by_2D from [12], trained on the Mapillary Vistas dataset. Differently from [12], we use the RGB colorized pointclouds for SK and NS instead of Lidar intensities, to reduce the domain gap with the MV dataset. We use Mask2Former models for both methods with the same parameters as IV-B. For Seg_3D_by_PC2D , we pretrain the 2D model on a PC2D dataset as in IV-B, and then fine-tune the model on the MV dataset for 10,000 iterations, while freezing the first block of the Swin backbone.
IV-E2 Results
Table IV reports the results as computed on SK for the traffic sign class. We see a significant improvement over the base method, with a 6.45 mIoU improvement. This shows that our method can be used to segment rare classes for which 3D annotations are not available.
@@ IV-E2 Results
| Seg_3D_by_2D | [12] | 11.85 | |
| Seg_3D_by_PC2D | (ours) | 18.3 | +6.45 |
IV-F Ablation Study
We conduct a complete ablation study to verify the importance of each component of our method. Results can be seen in Table V. The first stage (a) corresponds to seg_3D_by_2D from [12]. We add the PC2D dataset training is stage (b), the occlusion handling in stage (c), the View Generation Optimization in stage (d), the Multi-Modalities module in stage (e), and the full Seg_3D_by_PC2D is evaluated in stage (f). We evaluate the models in both SK10SK10 (in-domain) and NSSK10 (UDA).
We see that the addition of the PC2D dataset leads to the most significant improvement, followed by the occlusion handling and the multi-modalities module. Interestingly, the addition of the VGO module leads to an improvement only for the SK10SK10 setting, and not for UDA. This might be because using more diverse camera poses increases the risk of encountering a large domain shift within one camera pose type. It can also be noted that the addition of the RGB colorization only lead to a small improvement, which shows the method is applicable for datasets lacking RGB images. Overall, all components proposed in our method seems valuable for either supervised or UDA applications.
| PC2D | OCL | VGO | MMods | RGB | SK10 | NS | |
| (a) | 15 | 15 | |||||
| (b) | ✓ | 31.1 | 28.9 | ||||
| (c) | ✓ | ✓ | 42.0 | 34.4 | |||
| (d) | ✓ | ✓ | ✓ | 43.7 | 34.3 | ||
| (e) | ✓ | ✓ | ✓ | ✓ | 49.3 | 40.9 | |
| (f) | ✓ | ✓ | ✓ | ✓ | ✓ | 49.4 | 41.1 |
V Conclusion
We have presented a novel multi-view projection framework for 3D semantic segmentation that is especially well suited for UDA. After generating large-scale synthetic 2D datasets (PC2D) from aligned Lidar scenes to train an ensemble of 2D segmentation models, our method leverages occlusion-aware back-projection and voting to produce high-quality 3D labels without requiring any target features or annotations.
Extensive experiments demonstrate that our approach achieves state-of-the-art performance in UDA and can be used to segment ‘rare’ classes, with only 2D annotations for the class and 3D annotations for other classes are available.
While our approach thoroughly explored problem mentionned in Section II-B of what views to use for inference, and which models to infer with, interesting research could be done exploring problem of how to optimally use the numerous outputs from the ensemble of models on all views to assign final 3D labels. In particular, our simple logits accumulation and voting scheme does not use important information such as spatial proximity of the points or class size statistics. Information like this and other might be used to train a model at outputting the final 3D segmentation mask with more insight than a simple voting scheme.

### v3

@@ 4.1 Datasets and Metrics
Real-to-Real. SemanticKITTI [1] (SK) and nuScenes [2] (NS) are two well-known large-scale datasets for AD, both providing and annotating Lidar 3D point clouds. The procedure generally adopted in the domain adaptation literature [10, 11, 7, 24]
@@
is to train on the official training splits for both datasets, e.g. 700 scenes for nuScenes and sequences 00 to 07, 09 and 10 for SemanticKITTI, and evaluate on the official validation splits, e.g. 150 scenes for nuScenes and sequences 08 for SemanticKITTI. We follow this procedure for comparing with the state of the art. Of the 700 nuScenes scenes, we use the 70 last scenes for validation. Similarly, we use sequence 10 of SemanticKITTI for validation. For the rare classes experiment 4.6.2 and the ablation study 4.5, we use a subset of the first 1800 scans from sequence 08 of the SK dataset for evaluation, to reduce the computational cost.
@@
Simulation-to-Real. SynLidar [29] (SL) is a synthetic AD dataset captured with Unreal Engine 4. The sub-sampled version that is used in most of the literature [29, 23, 18, 30] is composed of 13 sequences, containing 19,865 scans with annotations for 32 classes. No official validation split is provided, therefore we follow [10] and use sequences 05 and 10 for validation. We do not evaluate on SynLidar, as it is only used as a source dataset.
Class mappings. We evaluate two Real-to-Real settings, NSSK and SKNS, and one Sim-to-Real setting, SLSK. To compare with the state-of-the-art, we follow [24, 18] and use 10 classes common to NS and SK. For the SLSK scenario, we use the official SK19 class mapping as in [23, 18]. To denote the different versions of the SemanticKITTI dataset, we use the following notations: SK10, SK19. The exact class mappings used can be found on github111https://github.com/andrewcaunes/ia4markings.
@@
Metrics. We measure the 3D semantic segmentation performance using the Intersection over Union (IoU) metric for all the experiments. We report both the per-class IoU and the mean IoU over all classes (mIoU). We compute all the metrics using code provided by [19], to ensure a fair comparison. For all UDA experiments, the evaluation is done with the 3D model after fine-tuning, to ensure similar conditions to the state-of-the-art methods, i.e. real-time inference on single scans. For the rare classes and ablation experiments, the evaluation is directly done on the pseudo-labels generated by the pipeline.
4.2 Implementation details
The most important hyper-parameters are presented in this section, but more details can be found on github††footnotemark: .
To begin with, we prepare the aligned scenes from SemanticKITTI and nuScenes. We use the already existing scenes from NS and accumulate scans 60 by 60 to create SK scenes. The intensities are normalized by scan, following [3].
@@ 4.2 Implementation details
NuScenes’ annotations are only available at 2Hz, while the scans are captured at 20Hz. To prepare for 2D segmentation label generation for PC2D datasets, we propagate the annotations to all points with a simple KNN method with to assign labels to each point in the dense scene.
@@
The PC2D datasets are generated with 400,000 samples for each modality and source dataset. The views are uniformly sampled over the training scenes. For inference, we use 250 images per camera pose type, thus 1,000 images per scene, per modality. This amounts to 4,000 images total per scene. All views are rendered at 1024512 resolution with a standard field of view. The inference time for each scene takes ~ 2 minutes on a NVIDIA RTX 3070 GPU. This is however only done once for pseudo-label generation, and there is no overhead introduced to the final fine-tuned model in both memory and time.
For the back-projection step, we crop the projection outside a depth range of [1m, 30m], and when occlusion is used, we set the occlusion margin to behind the depth encountered.
@@
4.3 Baselines
@@
To demonstrate the capabilities of our method, we compare it to the state-of-the-art methods in UDA for 3D semantic segmentation. We use the results reported by the authors for each method, while ensuring a fair comparison by using the same dataset splits, class mappings, metrics and model architectures. For Real-to-Real, in the NSSK10 setting, we find the best performing methods to date are Lidar-UDA [24], T-UDA [7], CoSMix [23] and SALUDA [18]. In the SK10NS setting, we compare with Lidar-UDA [24] and T-UDA [7], as CoSMix [23] and SALUDA [18] do not provide results for this setting. Note that T-UDA [7] also reports values for the manmade class, not reported by the other methods. We therefore do not include it and re-compute the mIoU accordingly. For Sim-to-Real, in the SLSK19 setting, we compare with SALUDA [18], CoSMix [23] and UniMix [30].
NSSK10 (% IoU) Car Bicycle Motorcycle Truck Other vehicle Pedestrian Driveable surf. Sidewalk Terrain Vegetation % mIoU Source-only 74.1 0.0 10.2 1.1 1.0 4.4 63.4 31.8 36.0 25.6 24.8 T-UDA [7] 93.0 0.0 11.4 3.4 47.0 15.7 83.3 54.4 67.9 83.9 46.0 Lidar-UDA [24] 86.2 0.0 13.9 9.3 3.1 16.5 65.7 6.1 54.1 85.7 34.1 CoSMix [23] 77.1 10.4 20.0 15.2 6.6 51.0 52.1 31.8 34.5 84.8 38.3 SALUDA [18] 89.8 13.2 26.2 15.3 7.0 37.6 79.0 50.4 55.0 88.3 46.2 Ours 90.1 0.0 28.4 17.7 18.4 37.1 80.9 61.0 70.2 90.1 49.4
SK10NS (% IoU) Car Bicycle Motorcycle Truck Other vehicle Pedestrian Driveable surf. Sidewalk Terrain Vegetation % mIoU Source-only 1.8 0.0 0.0 0.0 0.0 0.5 6.7 0.1 0.1 0.3 0.9 T-UDA [7] 74.2 0.5 40.3 21.8 0.2 0.4 87.8 45.8 46.1 70.3 38.7 Lidar-UDA [24] 73.5 0.9 15.9 0.9 25.7 40.8 87.4 42.3 47.9 83.2 41.8 Ours 73.3 0.5 32.0 50.7 0.0 32.1 90.3 55.4 58.2 87.0 47.9
4.4 Unsupervised Domain Adaptation Results
4.4.1 Real-to-Real setting
@@
Results for the NSSK10 setting can be found in Table 1. This UDA setting is particularly challenging as the source (NS) uses a 32-beam Lidar, less dense than the 64-beam Lidar used in the target (SK). Seg3DbyPC2D is able to outperform the state-of-the-art by mIoU points. Interestingly, our method reaches state-of-the-art performance even for dynamic classes such as car, motorcycle and truck. This demonstrates that the model learns to match ‘shadow effect’-like features with the corresponding classes.
@@
Results for the SK10NS setting can be found in Table 2. In this setting, Seg3DbyPC2D is even more dominant, with a mIoU advantage over state-of-the-art. The method results in particularly strong performance on static classes such as driveable surface, sidewalk, terrain and vegetation, which can be explained by the higher difficulty to segment dynamic classes displaying some ‘shadow effect’. For both tasks, we notice relatively lower performance on small classes, such as bicycle and pedestrian, which might be due to the projection errors caused by poor sensor calibration being more important for small objects.
SLSK19 (% IoU) Car Bicycle Motorcycle Truck Other vehicle Pedestrian Bicyclist Motorcyclist Road Parking Sidewalk Other ground Building Fence Vegetation Trunk Terrain Pole Traffic sign % mIoU Source-only 50.8 7.6 11.8 1.4 3.2 15.8 34.8 3.9 30.0 4.5 34.5 0 27.0 15.8 62.9 24.4 43.0 18.9 6.4 20.9 CoSMix [23] 75.1 6.8 29.4 27.1 11.1 22.1 25.0 24.7 79.3 14.9 46.7 0.1 53.4 13 67.7 31.4 32.1 37.9 13.4 32.2 SALUDA [18] 67.0 7.7 14.4 1.3 5.2 24.1 52.6 2.7 52.5 10.5 44.1 0.4 51.8 13.6 69.7 40.5 56.5 45.0 14.3 30.2 Unimix [30] 80.3 6.1 32.5 29.2 10.6 23.7 30 24.5 62.2 15.9 46.5 0.1 60.9 15.8 70.7 34.9 41.2 38.6 17.1 33.7 Ours 83.5 0.9 6 10.4 3.0 19.8 13.5 0.2 80.2 0.0 51.6 0 78.8 16.7 51.9 32.0 36.8 40.6 10.1 28.2
4.4.2 Simulation-to-Real setting
@@
Table 3 reports the results for the SLSK19 setting. Overall, our method achieves lower mIoU compared to prior works such as CoSMix [23], SALUDA [18], and UniMix [30]. However, it consistently performs strongly on several static classes, notably, road (80.2), sidewalk (51.6), and building (78.8), and some dynamic classes such as car (83.5), where it reaches state-of-the-art performance. This suggests that the projection-based pseudo-labeling is particularly effective on large, well-structured classes but less robust on fine-grained or rare classes such as bicycle, motorcycle, and motorcyclist. While the global mIoU is lower (28.2), the results highlight that our approach can still surpass prior methods on key static categories, underlining its strength for Sim-to-Real in some applications such as road infrastructure analysis.
4.5 Ablation Study
We conduct a complete ablation study to verify the importance of each component of our method. Results can be seen in Table 4. The first stage (a) corresponds to the base multi-view projection method that we followed from [3]. We add the PC2D dataset training in stage (b), the occlusion handling in stage (c), the Virtual Camera Placement in stage (d), and the full Seg3DbyPC2D with the Multi-Modalities module is evaluated in stage (e). We evaluate the models in both SK10SK10 (in-domain) and NSSK10 (UDA).
| PC2D | OCL | VCP | MMods | SK10 | NS | |||
| (a) | 15 | - | 15 | - | ||||
| (b) | ✓ | 31.1 | +16.1 | 28.9 | +13.9 | |||
| (c) | ✓ | ✓ | 42.0 | +10.9 | 34.4 | +5.5 | ||
| (d) | ✓ | ✓ | ✓ | 43.7 | +1.7 | 34.3 | -0.1 | |
| (e) | ✓ | ✓ | ✓ | ✓ | 49.3 | +5.6 | 40.9 | +6.6 |
We see that the addition of the PC2D dataset leads to the most significant improvement, followed by the occlusion handling and the multi-modalities module. Interestingly, the addition of the VCP module leads to an improvement only for the SK10SK10 setting, and not for UDA. This might be because using more diverse camera poses increases the risk of encountering a large domain shift within one camera pose type, e.g. the difference in Lidar point density might yield a larger difference in BEV type views than in road user type views, since the difference in point density becomes significant when seen from afar.
Overall, all components proposed in our method seems valuable for either supervised or UDA applications.
4.6 Application to rare classes
@@
A common real world scenario in 3D semantic segmentation is to have access to 3D annotations for some classes in a source domain, but not the desired target classes. However, 2D annotations are more generally available as they are less costly. For such a situation when 2D annotations of the target classes are available, [3] proposes a multi-view projection method with 2D models trained on real camera images, which then perform out-of-domain inference on virtual rendered views. We show that our method can help to significantly improve the performance of this approach by using the 3D annotations and our PC2D dataset generation pipeline to pretrain the 2D model in rendered view domain. The model is then fine-tuned on real camera images with annotations for the target classes. The pretraining on 3D annotations allows to significantly bridge the domain gap that exists at inference time by first exposing the model to rendered images.
4.6.1 Experimental setup
For this experiment, 3 datasets are used:
@@ 4.6.1 Experimental setup
A pretraining 3D source dataset, with 3D annotations for any classes (but not the target classes). We use nuScenes (NS) [2].
@@
A target 3D dataset. We use SemanticKITTI (SK) [1], as it conveniently provides 3D annotations for our target class, traffic sign, which will allow to perform a quantitative evaluation.
We compare our Seg3DbyPC2D with the baseline from [3], both trained on the rare classes dataset. Differently from [3] and from our base method in Sec. 3, we use the RGB colorized point clouds using camera images for SK and NS as the rendered views’ modality instead of Lidar intensities or other modalities. This reduces the domain gap with the real camera images of the rare classes dataset, but requires that camera data be available for the pretraining dataset. We use Mask2Former models for both methods with the same parameters as 4.2. For Seg3DbyPC2D , we pretrain the 2D model on a PC2D dataset, generated from the pretraining dataset, as in 4.2, and then fine-tune the model on the rare classes dataset for 10,000 iterations, while freezing the first block of the Swin backbone.
@@
| Baseline | [3] | 11.9 | |
| Seg3DbyPC2D | (ours) | 18.3 | +6.4 (+53%) |
4.6.2 Results
Table 5 reports the results as computed on SK for the traffic sign class. We see a significant improvement over the base method, with a mIoU improvement, corresponding to relative increase of . This shows that our method can be used to segment rare classes for which 3D annotations are not available, when some 2D annotations are available.
5 Conclusion
We have presented a new UDA method based on a novel multi-view projection framework for 3D semantic segmentation. After generating large-scale synthetic 2D datasets (PC2D) from aligned Lidar scenes in the source domain to train an ensemble of 2D segmentation models, our method leverages occlusion-aware back-projection and voting to produce high-quality 3D pseudo-labels for the target domain.
Extensive experiments demonstrate that our approach achieves state-of-the-art performance in Real-to-Real UDA and useful performance for large classes in Sim-to-Real UDA. We also demonstrated a potential application of our method to segment rare classes with no 3D annotations but 2D annotations available.
While our approach thoroughly explored problem mentionned in Section 2.2 of what views to use for inference, and which models to infer with, future research could be led to explore problem of how to optimally use the numerous outputs from the ensemble of models on all views to assign final 3D labels. In particular, our logits accumulation and voting scheme does not use important information such as spatial proximity of the points or class size statistics. Information like this could be used to train a model to output the final 3D segmentation mask with more insight than a simple voting scheme.